Estimate the beta of the following stock: market risk premium = 24.8 percent, RF = 5.8 percent, P0= $9.8, expected dividend at the end of the year = $2.30, P1= $12.30. Assume the market is in equilibrium.
Answer #1

Answer:

Required Return = (P1 + D1) / P0 - 1
Required Return = ($12.30 + $2.30) / $9.8 – 1
Required Return = 1.4898 – 1
Required Return = 0.4898

Required Return = Risk free rate + Beta * Market Risk Premium
0.4898 = 0.058 + Beta * 0.248
0.4898 – 0.058 = Beta * 0.248
0.4318 = Beta * 0.248
Beta = 0.4318 / 0.248
Beta = 1.74